OpenStudy (anonymous):

Find the six trig functions of the following: Theta=3pi/4

OpenStudy (rock_mit182):

@ForeverPurple \[\theta =\frac{ 3\pi }{ 4 }\]

OpenStudy (rock_mit182):

or 3(180°)/4 in degrees

OpenStudy (rock_mit182):

\[\sin \frac{ 3\pi }{ 4 } =\]

OpenStudy (rock_mit182):

\[\cos \frac{ 3\pi }{ 4 }\]

OpenStudy (rock_mit182):

\[\tan \frac{ 3\pi }{ 4 }=\]

OpenStudy (rock_mit182):

and so on . .

OpenStudy (rock_mit182):

use a calculator

OpenStudy (anonymous):

or the unit circle